# Distinguish the EMRI and B-EMRI system by gravitational waves

### v1

@@ 2 Numerical Kludge Waveforms
With all the results above, we are able to calculate the GW numerically with the NK method. In Figs.2, we exhibit the EMRI $h_{+}$ waveforms of single and binary stars for two cases $p=10M$ and $p=20M$. The orientation of the internal elliptic orbit plane is determined by the angles $\theta_{0}=\frac{\pi}{3},\phi_{0}=\frac{\pi}{6}$. The observer is supposed to be located at $r=D$ (100 Mpc in this case), $\Theta=\frac{\pi}{3}$, $\Phi=0$. From the figure we see that, the waveforms of the binary have the identical profile as that of the single star, but the waveforms of the binary have double spikes which are caused by the internal motion of the binary revolving around each other. We will analyze the waveforms in detail in the next section. In Fig.3, the orbits of one star in the binary system are displayed. It’s evident to see that distinct waveforms emanate from distinct orbits of motions.
Figure 7: Comparisons of amplitudes of GWs generated by pure external (first pannel) and that generated by pure internal motions (left three pannels). The mass of central SBH is $10^{6}M_{\odot}$, and the masses of the binary are $8M_{\odot}$ and $10M_{\odot}$. The initial external motion parameters are set to be $p(0)=10M,e(0)=0.3$, the internal motion parameters are fixed to be $\tilde{e}=0.2$ and $\tilde{p}=2M,0.5M,0.2M$.
@@ 3 Analysis of the waveforms
In this section, we compare the waveforms of binary stars with that of single star in detail. Through calculating the mismatch we analyze the degree of match between the two waveforms. The overlap between two signals is defined as
@@
$\tilde{x}^{*}(f)$ denote complex conjugate of $\tilde{x}(f)$, and $S_{h}(f)$ in (24) is the one-sided noise power spectral density of LISA [59], which is useful for surveying the types of sources that can be detected by the LISA mission.
Mismatch of two singles is defined as
@@
Figure 8: The $h_{+}$ waveform emanating from single (first panel) and binary compact objects (left panels) revolving around central SBH with mass $10^{8}M_{\odot}$. The masses of binary are set to be $8M_{\odot}$ and $10M_{\odot}$. The external orbit parameters are set to be $p(0)=10M,e(0)=0.3$. The internal motion parameters are set to be $\tilde{e}=0.2$ and $\tilde{p}=2M,M,0.5M,0.4M,0.2M$.
@@
where the SNR $\rho$ is defined as $\rho=\sqrt{\langle h,h\rangle}$ for a signal $h$. For LISA, the SNR is taken to be $\rho=20$[16].
Fig.4 displays the change of waveforms according to semi-latus rectum $\tilde{p}$ of binary internal elliptic orbit. We can see, compared with the waveform of single compact object (first panel in Fig.4) spikes appear in the waveforms of binary, and as $\tilde{p}$ decreases more spikes appear. In other words, the smaller $\tilde{p}$ is the larger the departure of binary waveforms from that of single star is. Thus it’s straightforward to guess the mismatch of the two types of waveforms increases as $\tilde{p}$ decreases, this is confirmed by the calculations of mismatch in frequency-domain in the following. In Fig.5, we compare the amplitudes of the GWs generated by pure external and that generated by pure internal motions. From the figure we see that, for the system composed of the central SBH with mass $10^{6}M_{\odot}$ and the binary with masses $1M_{\odot}$ and $1.1M_{\odot}$, if $p(0)$ is set to be $10M$, when the ratio of semi-latus rectums of internal and external orbits $\tilde{p}/p$ is about $1:50$, the amplitude of GWs generated by internal motions is comparable to that generated by external motions.
We fix the mass of central SBH to be $10^{6}M_{\odot}$ and increase the masses of binary to be $8M_{\odot}$ and $10M_{\odot}$, Fig.6 displays the waveforms of this B-EMRIs. Compared with Fig.4 we see that spikes also appear in the waveforms of the binary, and the amplitudes of GWs increase with the masses of the binary as expected. We compare the amplitudes of GWs generated by pure external and pure internal motions for this case as well, as can be seen in Fig.7, we get similar results as above, i.e., when the ratio of semi-latus rectums $\tilde{p}/p$ is about $1:50$, the amplitude of GWs generated by internal motions is comparable to that generated by external motions.
Since there are a large number of SBHs with mass larger than $10^{6}M_{\odot}$ in the universe, we would like to calculate the GWs emanating from the binary revolving around heavier SBHs. Fig.8 displays the GWs emanating from the B-EMRIs composed of the central SBH with mass $10^{8}M_{\odot}$ and the binary with masses $8M_{\odot}$ and $10M_{\odot}$. Similar to the above cases, due to period of binary internal motions decreases with $\tilde{p}$, more and more spikes appear as $\tilde{p}$ decreases. Now, for the initial parameter $p(0)=10M$ if the ratio of semi-latus rectums $\tilde{p}/p$ is about $3:10000$ (see Fig.9), the amplitude of GWs generated by internal motions is comparable to that generated by external motions.
To analyze the waveforms of single and binary compact objects more accurately we calculate the mismatch of the two involved waveforms. The first two panels in Fig.10 display the variation of mismatch with $\tilde{p}$, it shows that as $\tilde{p}$ decreases mismatch increases, which agrees with what we see in Fig.4. At first glance, one may expect the mismatch increases with $\tilde{p}$ since larger $\tilde{p}$ means greater departure from the orbits of single star. However, calculation shows as $\tilde{p}$ increases mismatch decreases instead. This is because the GWs are functions of derivatives of coordinates, the smaller $\tilde{p}$ is, the more greatly the coordinates of mass point change, and the more the departure of binary GW from that of single star is, and vice versa. We utilize the distinguishable criterion to judge whether the two involved waveforms are distinguishable. The dashed horizonal lines on the first two panels in Fig.10 denotes the indistinguishable bound corresponding to SNR $\rho=20$. It can be seen easily that, for small enough $\tilde{p}$ the GW signals of binary can be distinguished from that of single star, and that even if $\tilde{p}$ is not small enough such that the amplitude of GW generated by pure internal motions is small and can be treated as perturbations of GWs generated by pure external motions, the two waveforms can be distinguished. The third panel of Fig.10 displays the variation of the mismatch with $\theta_{0}$, it shows that, as $\theta_{0}$ increases mismatch first decreases and then increases, especially when $\theta_{0}=0$ mismatch takes maximum value, i.e., when the rotation axis of the binary is parallel to the rotation axis of the SBH, the waveforms of binary are easiest distinguished from that of the single star.
Figure 9: Comparisons of amplitudes of GWs generated by pure external (first pannel) and pure internal (left three pannels) motions of binary in Fig.8, the semi-latus rectum of internal orbits are set to be $\tilde{p}=0.2M,0.01M,0.003M$ respectively.
4 Discussion and Conclusion
@@
In this paper we study the GWs emanating from B-EMRIs which is composed of a central SBH and binary revolving around it. The SBH is supposed to be Kerr. We solve the GW equations by taking into account the mass quadrupole, mass octupole and current quadrupole moments of the source. For the radiation reaction we take the hybrid 2nd order post-Newtonian approximation which is applicable to generic inclined-eccentric orbits. We plot the GW waveforms of B-EMRIs and compare the waveforms of B-EMRIs with that of EMRIs of single star. The results show that, compared to the waveforms generated by single stellar-mass compact object the ones generated by binary compact objects exhibit double spikes or multi spikes. The spikes are caused by internal motions of the binary. The smaller $\tilde{p}$ is, the more and larger the spikes are. We compare the amplitudes of GWs generated by pure internal motions with that generated by pure external motions. Calculations show that for the B-EMRIs composed of central SBH with $M=10^{6}M_{\odot}$ and the binary with $m_{1}=M_{\odot},m_{2}=1.1M_{\odot}$ or $m_{1}=8M_{\odot},m_{2}=10M_{\odot}$, if the ratio of semi-latus rectums $\tilde{p}/p$ is about $1:50$, the amplitudes of GWs generated by pure internal motions is comparable to that generated by pure external motions. For the B-EMRIs composed of central SBH with $M=10^{8}M_{\odot}$ and the binary with $m_{1}=8M_{\odot},m_{2}=10M_{\odot}$, if the ratio $\tilde{p}/p$ is about $3:10000$, the amplitudes of GWs generated by pure internal motions is comparable to that generated by pure external motions.
We calculate the mismatch of two waveforms, the results show that, as $\tilde{p}$ decreases the mismatch increases, which agrees with the waveforms show us. This is a little counter-intuitive since it seems larger $\tilde{p}$ means greater departure from the orbits of single star. Actually, since GWs are functions of derivatives of coordinates, the larger $\tilde{p}$ is the smoother the orbit is, and the smaller the change rate of coordinates is, so larger $\tilde{p}$ corresponds to less departure of waveforms from that of single star. Applying the distinguishable criterion of two waveforms we find if $\tilde{p}$ is small enough the waveforms of binary are distinguishable from that of single star, even if $\tilde{p}$ is not too small and the GWs generated by pure internal motions can be viewed as perturbations of the GWs generated by pure external motions, the waveforms of B-EMRIs can be distinguishable from that of single star EMRIs. We also calculate the mismatch of two waveforms with different $\theta_{0}$ which is the angle between the surface of binary internal elliptic orbit and the horizonal surface, the results show that when $\theta_{0}=0$ mismatch takes maximum value, which means when the rotation axis of the binary is parallel to the rotation axis of the SBH, the waveforms of binary are in the most striking contrast from that of single star, and thus is the most beneficial position for observation to distinguish the GWs from EMRIs of single star and binary stars.
@@
Figure 10: Mismatch $\Delta$ as functions of $\tilde{p}$ or $\theta_{0}$. For the left panel the parameters are taken to be $p(0)=10M,e(0)=0.3,a=0.9,\tilde{e}=0.2,\theta_{0}=\frac{\pi}{3},m_{2}=1.1M_{\odot},m_{1}=M_{\odot}$, and $\tilde{p}=3M,4M,5M,5.5M,6M$ respectively. For the middle panel the parameters are fixed to be $p(0)=10M,e(0)=0.3,a=0.9,\tilde{e}=0.2,\theta_{0}=\frac{\pi}{3},m_{2}=10M_{\odot},m_{1}=8M_{\odot}$ and $\tilde{p}=0.2M,0.4M,0.5M,M,2M$ respectively. For the right panel the parameters are fixed as $p(0)=10M,e(0)=0.3,a=0.9,\tilde{p}=3M,\tilde{e}=0.2,m_{2}=1.1M_{\odot},m_{1}=M_{\odot}$, and $\theta_{0}=\frac{\pi}{12},\frac{\pi}{6},\frac{\pi}{4},\frac{\pi}{3},\frac{5\pi}{12},\frac{\pi}{2}$ respectively.

### v2

@@ 2 Numerical Kludge Waveforms
With all the results above, we are able to calculate the GW numerically with the NK method. In Figs.3-6, we exhibit the $h_{+}$ waveforms of EMRIs and B-EMRIs for different parameter selections. The orientation of the internal elliptic orbit plane is determined by the angles $\theta_{0}$ and $\phi_{0}$, $\phi_{0}$ is fixed to be $\frac{\pi}{6}$. The observer is supposed to be located at $r=D$, $\Theta=\frac{\pi}{3}$, $\Phi=0$. From the figure we see that, the waveforms of B-EMRIs have the identical profile as that of EMRIs, but the waveforms of B-EMRIs have oscillations which are caused by the internal motion of the binary. We will analyze the waveforms in detail in the next section.
2.3 Gravito-electromagnetic force
The internal orbit is calculated by Newtonian dynamics in Eq. 13. It is interesting and important to consider the higher order corrections e.g. the Gravito-electromagnetic force effect [60]. Unlike single free-falling compact object which always locates at the origin of the free-fall frame (FFF), every compact objects of the binary system will not locate at the origin of FFF due to the internal interaction of binary. According to equivalence principle only at the origin of FFF the spacetime is flat, if there is departure from the origin gravity will play a role. So for the binary we take into account the gravitational effects caused by the departure from the origin of FFF. For simplicity, we consider a circular orbit on the equatorial plane in the following. The natural choice of FFF is Riemann normal coordinates
$ds^{2}=-\left(1+R_{0i0j}x^{i}x^{j}\right)c^{2}d\tau^{2}-\frac{4}{3}R_{0jik}x^{j}x^{k}d\tau dx^{i}+\left(\delta_{ij}-\frac{1}{3}R_{ikjl}x^{k}x^{l}\right)dx^{i}dx^{j}.$ (29)
If $R_{ikjl}$, which is small compared to $R_{0i0j}$ and $R_{0jik}$, is neglected, the metric above can be rewritten as
$ds^{2}=-\left(1-2\frac{\Phi}{c^{2}}\right)c^{2}d\tau^{2}-\frac{4}{c}(\vec{A}\cdot d\vec{x})cd\tau+\delta_{ij}dx^{i}dx^{j},$ (30)
where
$\Phi$ $=-\frac{1}{2}R_{0i0j}x^{i}x^{j},$ (31)
$A_{i}$ $=\frac{1}{3}R_{0jik}x^{j}x^{k},$
are the scalar and vector potentials of GEM fields. Analogous to the electromagnetic fields, the GEM force can be introduced
$\vec{F}=-m\vec{E}-2m\frac{\vec{v}}{c}\times\vec{B},$ (32)
where
$E_{i}$ $=R_{0i0j}x^{j},$ (33)
$B_{i}$ $=-\frac{1}{2}\epsilon_{ijk}R_{0l}^{\;\;\;jk}x^{l}$
In FFF the equations of motion of the compact object in binary system can be wriiten as
$m_{i}\frac{d^{2}\vec{x}_{i}}{d\tau^{2}}=-m_{i}m_{j}\frac{\vec{x}_{i}-\vec{x}_{j}}{|\vec{x}_{i}-\vec{x}_{j}|^{3}}+\vec{F}_{i}$ (34)
where $i=1,2$ label the compact objects which constitute the binary.
Figure 7: GWs generated by B-EMRIs on the equatorial plane with and without taking into account gravito-electromagnetic force. The mass of central SBH is set to be $10^{8}M_{\odot}$, and the masses of the binary are set to be $8M_{\odot}$ and $10M_{\odot}$. The external orbit parameters are set to be $p(0)=10M,e(0)=0.3,a=0.9M$, the semi-major axis and eccentricity of internal orbit are set to be $\tilde{a}=10^{5}(m_{1}+m_{2}),\tilde{e}=0.2$ initially. The frequency of GWs is slightly increased when GEM force is in consideration. The physical interpretation is that the effect of general relativity tends to strengthen gravity compared to the Newtonian gravity.
Figure 8: Frequency spectrum of the GWs in Fig.3.
Although the equations of motion can be written in a simple form in FFF, it is not straightforward to calculate GEM force in FFF since Riemann tensor is conventionally derived in local non-rotating frame (LNRF). The transformations between the coordinates of LNRF ($X^{t},X^{r},X^{\theta},X^{\phi}$) and BL coordinates are given by
$dX^{t}$ $=(\Sigma\Delta/A)^{1/2}dt,$ (35)
$dX^{r}$ $=(\Sigma/\Delta)^{1/2}dr,$
$dX^{\theta}$ $=\Sigma^{1/2}d\theta,$
$dX^{\phi}$ $=-\frac{2Mar\sin\theta}{(\Sigma A)^{1/2}}+(A/\Sigma)^{1/2}\sin\theta d\phi,$
where $A=(r^{2}+a^{2})^{2}-a^{2}\Delta\sin^{2}\theta$. It’s shown Riemann tensor in LNRF takes very simple form[61, 62]. For circular orbit on the equatorial plane, the free-fall observer is moving in the $X^{\phi}$ direction with the speed[61]
$u=\frac{\pm M^{1/2}(r\mp 2aM^{1/2}r^{1/2}+a^{2})}{\Delta^{1/2}(r^{3/2}\pm aM^{1/2})}$ (36)
where the upper signs refer to prograde orbits and the lower ones refer to retrograde orbits. The LNRF and the local inertial frame (LIF) are connected by a Lorentz transformation
$\Lambda^{\mu}_{\mu^{\prime}}=\left(\begin{array}[]{cccc}\gamma&0&0&-\gamma\beta\\ 0&1&0&0\\ 0&0&1&0\\ -\gamma\beta&0&0&\gamma\\ \end{array}\right),$ (37)
where $\beta=u/c,\gamma=\sqrt{1-\beta^{2}}$. Through the Lorentz transformation (37) Riemann tensor in LNRF can be transformed to the ones in LIF $R_{\mu^{\prime}\nu^{\prime}\rho^{\prime}\sigma^{\prime}}=\Lambda^{\mu}_{\mu^{\prime}}\Lambda^{\nu}_{\nu^{\prime}}\Lambda^{\rho}_{\rho^{\prime}}\Lambda^{\sigma}_{\sigma^{\prime}}R_{\mu\nu\rho\sigma}$. If a free gyro placed at the origin of FFF, its spin axis is fixed in FFF, but in general the gyro will precess relative to the LNRF and LIF, so there exist rotation between LIF and FFF, the angular velocity is $\omega=-\sqrt{M/r^{3}}$[63]. The tensors in LIF and the that in FFF are connected through a rotation transformation. The components of GEM force in FFF can be given in terms of that in LIF as
$F_{x}$ $=F_{r^{\prime}}\cos(\omega\tau)+F_{\phi^{\prime}}\sin(\omega\tau),$ (38)
$F_{y}$ $=F_{\theta^{\prime}},$
$F_{z}$ $=-F_{r^{\prime}}\sin(\omega\tau)+F_{\phi^{\prime}}\cos(\omega\tau).$
Substituting the GEM force (38) into the equations of motion (34) and solving the equations allows to obtain the positions of compact objects in FFF. Considering the orbit of origin of FFF is geodesic in external SBH spacetime, it’s straightforward to obtain the positions of compact objects of the binary in SBH spacetime, based on this one is able to calculate the GWs with the NK method. In Fig.7 we present the waveform of GW with and without taking into account GEM force.
@@ 3 Analysis of the waveforms
In this section, we compare the waveforms of B-EMRIs with that of EMRIs in detail. Furthermore, we investigate the modification of the waveform by considering the Gravito-electromagnetic (GEM) force effect and show that the dominated effect of GEM is a surplus phase shift. Fig.3 displays the $h_{+}$ waveforms of EMRIs and B-EMRIs with different masses. One sees that the profile of GWs of EMRIs and B-EMRIs are similar, but the details are distinct. For B-EMRIs, the gravitational waveforms possess small fluctuations. This is caused by the internal motions of the binary. We see, as mass increases, both the amplitudes and the frequency of the fluctuations increase. Fig.4 displays the the $h_{+}$ waveforms of EMRIs and B-EMRIs with different semi-latus rectum $\tilde{p}$ of the internal orbit. The figure shows that the smaller $\tilde{p}$ is, the larger the amplitude and the frequency of the fluctuation are, and vice versa. Since smaller $\tilde{p}$ corresponds to small cycle or higher frequency, amplitude of the fluctuation increases as $\tilde{p}$ decreases is expected.
Fig.5 displays the comparisons of the $h_{+}$ waveforms between EMRIs and B-EMRIs with different $\theta_{0}$, which shows that, as $\theta_{0}$ increases the amplitude of fluctuation first decreases and then increases, especially when $\theta_{0}=0$ or $\theta_{0}=\pi$ the amplitude of fluctuation takes maximum value, i.e., when the rotation axis of the binary is parallel or anti-parallel to the rotation axis of the SBH, the waveforms of B-EMRIs and EMRIs take the most sharp distinction.
In Fig.6 the $h_{+}$ waveforms of EMRIs and B-EMRIs with SBH weighted $M=10^{8}M_{\odot}$ are displayed. One sees that, since the central SBH is much heavier in this case, the amplitude of fluctuation generated by binary composed of compact objects weighted $8M_{\odot}$ and $10M_{\odot}$ is very tiny. This is because for heavier BH with parameter selection $p(0)=10M$ the cycle is much longer and frequency is much smaller. As displayed in Fig.3, like the $M=10^{6}M_{\odot}$ case as mass increases the amplitude of fluctuation increases.
Figure 9: Frequency spectrum of the GWs in Fig.4.
To compare the waveforms of B-EMRIs and EMRIs we analyse the waveforms in frequency domain. In Fig.8 we present the frequency spectrum of the waveforms in Fig.3. One sees that, in low frequency region the spectrum of B-EMRIs coincide with that of EMRIs. In high frequency region, only the B-EMRIs have non-vanishing amplitudes of gravitational signals. Further, as mass increases, the amplitudes of signals increase, and the more and more higher-frequency signals emerge. This result agrees with the waveforms in Fig.3, where it shows as mass increases the amplitude and frequency of the fluctuations become larger. This is because, on the one hand, the amplitude of GW is proportional to the mass of the source, on the other hand, as mass increases, the cycle decreases and the frequency increases. From the GW expression (15) we know, as frequency increases the amplitude of GW increases as well.
Figure 10: Frequency spectrum of the GWs in Fig.5. In high frequency region the spectrum of EMRIs is definitely different from the spectrum of B-EMRIs. We can distinguish B-EMRIs from EMRIs without question in frequency domain by using the data from space-based GW detectors.
In Fig.9, we give the frequency spectrum of the GWs in Fig.4. We can see, as $\tilde{p}$ decreases, the amplitude of GW signal increases, and more and more high-frequency signals emerge, and vice versa. This agrees with what the waveform shows. Since the smaller $\tilde{p}$ is, the shorter the cycle is, and the higher the frequency is, so the amplitude of the signal increases. In Fig.10 we give the frequency spectrum of the waveforms in Fig.5 with distinct $\theta_{0}$. We can see, for distinct $\theta_{0}$, the frequency range is not significantly different. The amplitude of the signals with different frequency first decreases and then increases. For $\theta_{0}=0$ or $\theta_{0}=\pi$, the amplitude of the signals takes the maximum value, which agrees with the waveforms in Fig.5.
When GEM force is taken into account, phase shift between the GW waveforms of B-EMRIs and that of EMRIs emerges, as shown in Fig.7. It’s sensitive and valuable to calculate mismatch of the two waveforms to quantify the distinction between them. The inner product of two signals is defined as
@@
Here * denotes the complex conjugate of a quantity, and $S_{h}(f)$ in (41) is the one-sided noise power spectral density of LISA [59], which is significant for surveying the types of sources that can be detected by the LISA mission.
With the inner product defined in (39), the overlap of two signals is given by
$\mathcal{O}=\frac{\langle x,h\rangle}{\sqrt{\langle x,x\rangle}\sqrt{\langle h,h\rangle}}.$ (41)
Mismatch of two GW signals is defined as
@@
where the SNR $\rho$ is defined as $\rho=\sqrt{\langle h,h\rangle}$ for a signal $h$. For LISA, the SNR is taken to be $\rho=20$[16]. Calculation shows for the two waveforms in the first panel of Fig.7 $\Delta=0.003921$, for the two waveforms in the second panel of Fig.7 $\Delta=0.03922$, which means when taking into account GEM force the waveforms of B-EMRIs are distinguishable from the waveforms of B-EMRIs that does not take into account GEM force, and are distinguishable from the waveforms of EMRIs as well.
4 Conclusion and Discussion
In this paper we study the GWs emanated from B-EMRIs which is composed of a central SBH and binary revolving around it. B-EMRIs are quite popular in the universe, and remarkable objects for GW detectors of next generation. We demonstrate that the waveform of B-EMRIs is distinctly different from that of EMRIs in frequency domain, and show that the mismatch can be effectively probed by the space-based GW detectors such as LISA, Taiji, and Tianqin. Our results will improve template construction for B-EMRIs and shed light on detection of GWs from B-EMRIs.
@@ 4 Conclusion and Discussion
We solve the GW equations by taking into account the mass quadrupole, mass octupole and current quadrupole moments of the source. For the radiation reaction we take the hybrid 2nd order post-Newtonian approximation which is applicable to generic inclined-eccentric orbits. We plot the GW waveforms of B-EMRIs and compare the waveforms of B-EMRIs with that of EMRIs of single star. The results show that, the waveforms generated by single stellar-mass compact object share the same profile with the ones generated by binary compact objects, but the details of the waveforms of B-EMRIs exhibit fluctuations. The fluctuations are caused by internal motions of the binary. As masses of the compact objects of binary increase, the amplitudes and frequencies of the fluctuations increase. This is because, on the one hand, the amplitude of GW is proportional to the mass of the source. On the other hand, the increase of mass of gravitational source cause the decrease of cycle, so the frequency increases, which increases the amplitude further. We fix other parameters and only adjust the semi-latus rectum of internal orbits $\tilde{p}$, the result shows that, as $\tilde{p}$ decreases, the amplitudes and frequencies of the fluctuations increase. This is because, as $\tilde{p}$ decreases, the cycle of internal motion decreases, the frequency increases, which causes the increase of amplitudes of GWs. We fix other parameters and adjust the angle $\theta_{0}$ between axes $z$ and $z_{0}$, the results show that, when $\theta_{0}=0$ or $\pi$ the amplitudes of GW fluctuations caused by binary internal motions are larger than the cases when $\theta_{0}$ takes other values, which means when $\theta_{0}=0$ or $\pi$ the GWs of B-EMRIs are most easily distinguished from that of EMRIs.
For the binary on the equatorial plane of SBH, we further take in account GEM force and calculate the GWs of B-EMRIs. By comparing with both the waveforms of B-EMRIs that does not take into account GEM force and the waveforms of EMRIs, it’s easy to find there is phase shift between the related two waveforms.
To compare the GWs of EMRIs and B-EMRIs further we perform the frequency spectrum analysis. The results show that, at low-frequency region, the frequency spectrum of B-EMRIs coincide with that of EMRIs. At high-frequency region, the frequency spectrum of B-EMRIs and EMRIs are different. As masses of the binary increase, more high-frequency signals emerge, and the amplitudes of the GW signals increase, which agrees with the waveforms show us. We give the frequency spectrum of GWs generated by binary with distinct semi-latus rectum of internal orbits $\tilde{p}$, the results show that, as $\tilde{p}$ decreases, more high-frequency signals emerge, and the amplitude of the GW signals increase, since smaller $\tilde{p}$ corresponds to shorter cycle and higher-frequency orbit motions, higher-frequency orbit motions generate GWs with higher amplitudes. If we adjust the angle $\theta_{0}$ between axes $z$ and $z_{0}$, frequency spectrum show that when $\theta_{0}=0$ or $\pi$ the amplitudes of GWs are larger than that of GWs when $\theta_{0}$ takes other values, this agrees with what the waveforms show us as well. So the GW signals of B-EMRIs can be clearly distinguished from that of EMRIs through spectrum analysis, the GW signals generated by binary with larger mass, closer distance are more easily to be distinguished from that generated by single compact object. Moreover, when rotation axis of the binary is parallel or anti-parallel to that of the central SBH, the GW signals of B-EMRIs are more easily to be distinguished from that of EMRIs. When taking into account GEM force, we analyze the waveforms by calculating mismatch, calculation shows the GW waveforms of B-EMRIs are distinguishable from the waveforms of B-EMRIs that does not take into account GEM force and from the waveforms of EMRIs.
Figure 11: Frequency spectrum of the GWs in Fig.6.